Find the joint pmf of two independent poissonally distributed random variables X and Y where

μ1 = μx and μ2 = μy. Then evaluate P(X + Y = 5).


1
Expert's answer
2021-11-25T13:38:30-0500

 pmf of X+Y:

"f=P(X+Y=K)=\frac{(\mu_x+\mu_y)^ke^{-\mu_x-\mu_y}}{k!}"


"P(X+Y=5)=\frac{(\mu_x+\mu_y)^5e^{-\mu_x-\mu_y}}{5!}"